OpenStudy (anonymous):

If 1/4 of an even positive number ad 5/6 of the next larger even number have a sum of 32, what is the average of the two numbers?

OpenStudy (anonymous):

29

OpenStudy (anonymous):

29

OpenStudy (anonymous):

Any even number can be represented as 2x. thus, (2x)/4 + 5(2x + 2)/6 = 32 simplifying... x/2 + 5x/3 + 5/3 = 32 13x/6 = 182/6 13x = 182 x = 14 Your first even number is therefore 2x = 28 the next larger even number is (2x + 2) = 30 the average of these is 29.